A group of students at Old Sawash U. who are well trained in statistics, makes a claim regarding their schools superiority.
Answer #1

Sample

n = 200

Mean = 520

Std Dev = 100

Std Error (SE) = 520/1001/2 = 52

Alpha = 0.05

Null and Alternate Hypothesis

H0: µ = 500

Ha: µ > 500

This is a one-tailed test

Test Statistic

t = (520-500)/52 = 0.385

tCritical = 1.653

p-value = TDIST(0.385,200-1,1) = 0.350

Result

Since the p-value is greater than 0.05, we fail to reject the null hypothesis.